# Stopping rules for Monte Carlo methods of martingale difference type

### v1

@@ 4 The stopping rules
Plainly speaking, the three assumptions here describe that over every single batch, its sample is neither completely degenerate nor exploding in variance, as well as the batch variance stays in a common order in the long run. It is worth noting that Theorem 2.4 (ii) holds true under the second assumption ($\sup_{t\in\mathbb{N}}v^{2}_{0}(t)<+\infty$) here, since the condition ($\limsup_{t\to+\infty}\sum_{k\in M(t)}(k-m(t-1))^{-2}{\rm Var}_{0}(X_{k})<+\infty$) in Theorem 2.4 (ii) is weaker. For instance, with ${\rm Var}_{0}(X_{k})\sim(k-m(t-1))^{q}$ for any $q\in(0,1)$, we have $v^{2}_{0}(t)\to+\infty$ while the condition $\limsup_{t\to+\infty}\sum_{k\in M(t)}(k-m(t-1))^{-2}{\rm Var}_{0}(X_{k})<+\infty$ holds true. Hence, the convergence $\lim_{t\to+\infty}\mu(t)=\mu$ remains true $\mathbb{P}_{0}$-$a.s.$, in what follows. The third assumption ($v^{2}_{0}(t+1)/|M(t+1)|\sim v^{2}_{0}(t)/|M(t)|$) does not look intuitive but is rather technical for asymptotic validation of the proposed stopping rules. Here, we present a standard form of the martingale central limit theorem [25] in our notation.
@@ 4.1 The stopping rule based on theoretical batch variances
$\tau^{0}_{\epsilon,\delta}:=\mathop{\rm argmin}_{t\in\mathbb{N}}\left\{\mathbb{P}_{0}(|\mu(t)-\mu|>\epsilon)\leq\delta\right\},$ (4.3)
with precision $\epsilon\in(0,+\infty)$ and error probability $\delta\in(0,1)$ specified in subscript for asymptotic justification, unlike in Section 3. As is obvious from its expression, the stopping batch $\tau_{\epsilon,\delta}^{0}$ is $\mathcal{F}_{0}$-measurable, that is, deterministic in effect, as well as is far from implementable because the law of $|\mu(t)-\mu|$ is unknown. Still, at this initiation phase, we justify its “asymptotic” relevance in a similar manner to [12], as if the precision $\epsilon$ and the error probability $\delta$ could be adjusted after the sequence $\{X_{k}\}_{k\in\mathbb{N}}$ of infinite length had been generated.
@@
$\tau^{0}_{\epsilon,\delta}$ $=\mathop{\rm argmin}_{t\in\mathbb{N}}\left\{\mathbb{P}_{0}(|\mu(t)-\mu|>\epsilon)\leq\delta\right\}$
$\sim\mathop{\rm argmin}_{t\in\mathbb{N}}\left\{2\left(1-\Phi\left(\frac{\epsilon\sqrt{|M(t)|}}{v_{0}(t)}\right)\right)\leq\delta\right\},$ (4.4)
@@
$\quad=\mathbb{P}_{m(\tau^{0}_{\epsilon,\delta})}\left(\sqrt{|M(\tau^{0}_{\epsilon,\delta})|}\frac{|\mu_{\star}(\tau^{0}_{\epsilon,\delta})-\mu|}{v(\tau^{0}_{\epsilon,\delta})}>\sqrt{|M(\tau^{0}_{\epsilon,\delta})|}\frac{\epsilon}{v(\tau^{0}_{\epsilon,\delta})}\right)$
$\quad\approx 2\left(1-\Phi(\epsilon\sqrt{|M(\tau^{0}_{\epsilon,\delta})|}/v(\tau^{0}_{\epsilon,\delta}))\right),$ (4.5)
provided that the value of the stopping batch $\tau_{\epsilon,\delta}^{0}$ is luckily far away. We stress that the last approximation in (4.5) is valid because the resampled batch $\{X_{k}^{\star}\}_{k\in M(\tau)}$ is conditionally independent of the stopping batch $\tau_{\epsilon,\delta}^{0}$ on the stopping-time $\sigma$-field $\mathcal{F}_{m(\tau_{\epsilon,\delta}^{0})}$.
@@ 4.2 The stopping rule based on implementable batch variances
As is now obvious from its expression, the crucial drawback of the approximate stopping rule (4.5) is the unavailability of the theoretical batch standard deviation $\{v_{0}(t)\}_{t\in\mathbb{N}}$. In what follows, we aim to replace the theoretical batch standard deviation in the approximate criterion (4.4) with a more implementable sequence $\{v(t)\}_{t\in\mathbb{N}}$ and make the resulting criterion more precise for a better application of the following martingale central limit theorem [27, Chapter VIII.1], which is a direct consequence of Theorem 4.2 owning to the continuous mapping theorem.
@@
In the literature, a variety of sufficient conditions have been provided in different forms for the weak convergence (4.6) to hold true, which we do not pursue as those are not directly relevant in our context. For detail, we refer the reader to, for instance, [16, Theorem 2.5]. Let us stress that the condition $v^{2}(t)/v^{2}_{0}(t)\stackrel{{\scriptstyle\mathbb{P}_{0}}}{{\to}}1$ implies that the sequence $\{v(t)\}_{t\in\mathbb{N}}$ does not vanish in the limit, due to Assumption 4.1. From a practical point of view, this is a reasonable property, since the sample $\{X_{k}\}_{k\in\mathbb{N}}$ itself (not the sample average) satisfying $v^{2}(t)\to 0$ would otherwise necessarily converge to the constant $\mu$ of interest, for which a stopping rule would no longer be of particular use.
@@
$\tau_{\epsilon,\delta}:=\mathop{\rm argmin}_{t\in\mathbb{N}}\left\{2\left(1-\Phi\left(\frac{\epsilon\sqrt{|M(t)|}}{v(t)+a(t)}\right)\right)\leq\delta\right\},$ (4.7)
@@
Let us emphasize that the sequence $\{v(t)\}_{t\in\mathbb{N}}$ in (4.7) cannot be built on the resampled sequence $\{X_{k}^{\star}\}_{k\in M(\tau_{\epsilon,\delta})}$, because $\tau_{\epsilon,\delta}$ needs to be an $(\mathcal{F}_{m(t)})_{t\in\mathbb{N}_{0}}$-stopping time, generated by the original sequence $\{X_{k}\}_{k\in\mathbb{N}}$ alone. On the basis of the stopping time (4.7), we construct a stopping rule as follows.
@@
In a similar manner to [3], the “asymptotic” relevance of the criterion (4.7) can be justified in a way that the precision $\epsilon$ and the error probability $\delta$ are adjusted after the sequence $\{X_{k}\}_{k\in\mathbb{N}}$ of infinite length (which is certainly impossible to keep in hand) has been generated. We note that the almost sure convergence $v^{2}(t)/v^{2}_{0}(t)\to 1$ can be weakened to the weak convergence $\{v^{2}(\lfloor t/\epsilon\rfloor)/v^{2}_{0}(\lfloor t/\epsilon\rfloor):\,t\in(0,+\infty)\}\to 1$ in $\mathbb{D}((0,+\infty);\mathbb{R})$ as $\epsilon\to 0+$ for a version of Proposition 4.5 based on the functional weak law of large numbers, while it cannot be weakened further to $v^{2}(t)/v^{2}_{0}(t)\stackrel{{\scriptstyle\mathbb{P}_{0}}}{{\to}}1$ since then the corresponding version of Part (ii) would not hold anymore (see, for instance, [13]).
@@
In relation to Proposition 4.5 (ii), we stress that $\mathbb{P}_{0}(|\mu(\tau_{\epsilon,\delta})-\mu|>\epsilon)\not\to\delta$ in general, as the random variable $\mu(\tau_{\epsilon,\delta})$ is not identical in law to the empirical batch mean $\mu(t)$ at a usual deterministic batch $t$ in the presence of the stopping time $\tau_{\epsilon,\delta}$. Conversely, the conditional standard deviation $s(\tau_{\epsilon,\delta})$ at the stopping batch $\tau_{\epsilon,\delta}$ cannot be replaced by its resampled version, that is, $v_{\star}^{2}(\tau_{\epsilon,\delta})=|M(\tau_{\epsilon,\delta})|^{-1}\sum_{k\in M(\tau_{\epsilon,\delta})}{\rm Var}_{k-1}^{\star}(X^{\star}_{k})$, since the resampled batch $\{X^{\star}_{k}\}_{k\in M(\tau)}$ is conditionally independent of the stopping batch $\tau_{\epsilon,\delta}$ on the stopping-time $\sigma$-field $\mathcal{F}_{m(\tau_{\epsilon,\delta})}$.
With Proposition 4.5 being presented, of practical interest is the non-asymptotic regime [1, 17] in which one generates a segment $\{X_{k}\}_{k\in M(t)}$ of the sequence, one batch at a time forward until, and only until, the criterion (4.7) is satisfied, with precision $\epsilon$ and error probability $\delta$ fixed.
@@ 4.2.1 Conditional batch variances
and $v_{1}(t):=\left(v_{1}^{2}(t)\right)^{1/2}$ for $t\in\mathbb{N}$, where we have applied the base assumption (2.1) for the evaluation of the conditional variance. The conditional batch variance here is expected to behave in a somewhat similar manner to the theoretical one in the sense that it is unbiased for the theoretical batch variance (4.1), that is, $\mathbb{E}_{0}[v_{1}^{2}(t)]=v^{2}_{0}(t)$ for all $t\in\mathbb{N}$. Based on the availability of the sample $\{X_{k}\}_{k\in M(t)}$ alone, the conditional batch variance and standard deviation (4.8) are, in principle, more implementable than the theoretical ones (4.1), whereas the reality is that neither is generally implementable, with a few exceptions (such as the one examined in Section 5.1, where the conditional one (4.8) happens to be available in an implementable form).
@@
$\tau^{a}_{\epsilon,\delta}:=\mathop{\rm argmin}_{t\in\mathbb{N}}\left\{2\left(1-\Phi\left(\frac{\epsilon\sqrt{|M(t)|}}{v_{1}(t)+a(t)}\right)\right)\leq\delta\right\},$ (4.9)
@@ 4.2.2 Empirical batch variances
$\tau^{b}_{\epsilon,\delta}:=\mathop{\rm argmin}_{t\in\mathbb{N}}\left\{2\left(1-\Phi\left(\frac{\epsilon\sqrt{|M(t)|}}{v_{2}(t)+a(t)}\right)\right)\leq\delta\right\},$ (4.11)
@@ 5 Numerical illustrations
In accordance with [1], we consider a stopping rule with its reliability $R_{\epsilon,\delta}(\omega)$ strictly less than $1$ unreliable. In addition, the effectiveness of a stopping rule improves as its complexity $\text{CM}_{\epsilon,\delta}(\omega)$ decreases at the specified precision and error probability. We note that the parameter $\ell$ in the complexity $\text{CM}_{\epsilon,\delta}(\omega)$ controls how severely we penalize large stopping times and is suppressed from its notation because we fix $\ell=5$ throughout. To ensure sufficient realizations, we set $n=5\times 10^{3}$ above.
@@ 5.1 ARCH(1)
We begin our numerical illustrations with an ARCH(1) model, which provides an ideal test case for our stopping rules due to its non-trivial dependence structure and heavy-tailed behavior, as well as the straightforward implementation of the theoretical and conditional batch variances. This setting poses natural challenges for sequential sampling methods, as the underlying dynamics combine temporal dependence through the ARCH structure with potentially extreme observations from the $t$ distribution.
Now, consider an ARCH(1) sequence $\{X_{k}\}_{k\in\mathbb{N}}$, which is generated recursively by $X_{k}=(\beta+\alpha X_{k-1}^{2})^{1/2}V_{k}$, with $X_{0}=0$, $\alpha\in(0,1)$ and $\beta\in(0,+\infty)$. Here, $\{V_{k}\}_{k\in\mathbb{N}}$ denotes a sequence of iid random variables following a scaled $t$-distribution with $n$ degrees of freedom ($n\in\{5,6,\cdots\}$), normalized to have unit variance, with its probability density function given as
$\frac{d}{dx}\mathbb{P}(V_{1}\leq x)=\frac{\Gamma((n+1)/2)}{\sqrt{(n-2)\pi}\Gamma(n/2)}\left(1+\frac{x^{2}}{n-2}\right)^{-(n+1)/2},\quad x\in\mathbb{R}.$
@@
In what follows, we fix $\alpha=0.03$, $\beta=0.3$, and $n=6$ degrees of freedom. The batch structure is set as $M(t)=\{(t-1)^{5}+1,\cdots,t^{5}\}$, ensuring sufficient sample growth while maintaining theoretical requirements.
Figure 2: Typical realization of theoretical batch variance $v_{0}^{2}(t)$ (solid blue), conditional batch variance $v_{1}^{2}(t)$ (dashed red), and empirical batch variance $v_{2}^{2}(t)$ (dotted green) of the ARCH(1) process under consideration.
@@
As we observe from Figure 2, after some initial fluctuations due to smaller batch sizes, the theoretical, conditional, and empirical variances exhibit similar behavior, with all those measures tracking each other closely. This validates the use of the empirical variance $v_{2}^{2}(t)$ in the stopping rule, as it succeeds in approximating the theoretical variance, not to mention the conditional variance, in the long run.
@@
We implement $\tau^{b}_{\epsilon,\delta}$ based on the empirical batch variance $v_{2}^{2}(t)$ as of (5.1), as it can be readily available unlike the theoretical and conditional variances. Here, we set $a(t)=1/t$, which helps to compensate the severe underestimation of the empirical variance $v_{2}^{2}(t)$ during early batches, such as $v_{2}^{2}(1)\approx 0$ observed in Figure 2. The choice of $a(t)=1/t$ satisfies its theoretical requirement of monotonic vanishing and reduces the risk of premature stopping.
@@
We assess its performance by examining reliability and computational complexity ((iv)) across various precision ($\epsilon$) and error probability ($\delta$) values. For each $(\epsilon_{k},\delta_{k})$ in the specified grid, we perform independent simulations to compute $\tau^{b}_{\epsilon_{k},\delta_{k}}(\omega_{j})$, generate new batches for the resampled mean $\mu_{\star}(\tau_{\epsilon_{k},\delta_{k}})(\omega_{j})$, and evaluate success through $\mathbbm{1}(|\mu_{\star}(\tau_{\epsilon_{k},\delta_{k}})(\omega_{j})-\mu|\leq\epsilon_{k})$.
Mean Min Max
Reliability ($R_{\epsilon,\delta}$) 1.0225 1.0010 1.1111
Complexity ($\rm CM_{\epsilon,\delta}$) 1,601,368.29 456.92 16,417,500.00
Table 1: Performance metrics
In Figure 3 and Table 1 above, we present our comprehensive numerical findings on the performance of the proposed stopping rule. The reliability plot (the left plot of Figure 3) demonstrates that our stopping rules consistently achieve the reliability $R_{\epsilon,\delta}$ staying strictly greater than $1$ across all parameter combinations. As confirmed in Table 1, reliability values range from 1.001 to 1.111, with a mean of 1.0225, satisfying the reliability criterion set up in [1]. The reliability increases systematically with larger values of $\epsilon$ and $\delta$, reflecting more conservative behavior in less demanding scenarios while maintaining reliability even in the most challenging case with $\epsilon=\delta=10^{-3}$. The complexity heat map (the right plot of Figure 3) reveals a pattern of computational cost across the parameter space. The logarithmic scale clearly shows how the complexity $\rm CM_{\epsilon,\delta}$ increases as both $\epsilon$ and $\delta$ decrease. The computational burden varies substantially, as detailed in Table 1, ranging from $\rm CM_{\epsilon,\delta}\approx 4.57\times 10^{2}$ for larger values of $\epsilon$ and $\delta$ to $\rm CM_{\epsilon,\delta}\approx 1.64\times 10^{7}$ for $\epsilon=\delta=10^{-3}$, with a mean of approximately 1.6 million samples. This scaling aligns with our theoretical predictions and provides practical guidance for computational requirements. Smooth transitions in both reliability and complexity across the parameter space indicate the stability and robustness of the proposed stopping rule. The consistent achievement of the requirement $R_{\epsilon,\delta}>1$, coupled with theoretically predicted complexity scaling, provides strong empirical validation of our methodology, particularly notable given the challenging features of temporal dependence and heavy-tailed distributions in the ARCH(1) setting.

### v2

@@ 4 The stopping rules
Plainly speaking, the three assumptions here describe that over every single batch, its sample is neither completely degenerate nor exploding in variance, as well as the batch variance stays in a common order in the long run. It is worth noting that Theorem 2.4 (ii) holds true under the second assumption ($\sup_{t\in\mathbb{N}}v^{2}_{0}(t)<+\infty$) here, since the condition ($\limsup_{t\to+\infty}\sum_{k\in M(t)}(k-m(t-1))^{-2}{\rm Var}_{0}(X_{k})<+\infty$) in Theorem 2.4 (ii) is weaker. For instance, with ${\rm Var}_{0}(X_{k})\sim(k-m(t-1))^{q}$ for any $q\in(0,1)$, we have $v^{2}_{0}(t)\to+\infty$ while the condition $\limsup_{t\to+\infty}\sum_{k\in M(t)}(k-m(t-1))^{-2}{\rm Var}_{0}(X_{k})<+\infty$ holds true. Hence, the convergence $\lim_{t\to+\infty}\mu(t)=\mu$ remains true $\mathbb{P}_{0}$-$a.s.$, in what follows. The third assumption ($v^{2}_{0}(t+1)/|M(t+1)|\sim v^{2}_{0}(t)/|M(t)|$) does not look intuitive but is rather technical for asymptotic validation of the proposed stopping rules. Here, we present a standard form of the martingale central limit theorem [martingalebook2022] in our notation.
@@ 4.1 The stopping rule based on theoretical batch variances
$\tau^{0}_{\epsilon,\delta}:=\min\left\{t\in\mathbb{N}:\,\mathbb{P}_{0}(|\mu(t)-\mu|>\epsilon)\leq\delta\right\},$ (4.3)
with precision $\epsilon\in(0,+\infty)$ and error probability $\delta\in(0,1)$ specified in subscript for asymptotic justification, unlike in Section 3. As is obvious from its expression, the stopping batch $\tau_{\epsilon,\delta}^{0}$ is $\mathcal{F}_{0}$-measurable, that is, deterministic in effect, as well as is far from implementable because the law of $|\mu(t)-\mu|$ is unknown. Still, at this initiation phase, we justify its “asymptotic” relevance in a similar manner to [PW1992], as if the precision $\epsilon$ and the error probability $\delta$ could be adjusted after the sequence $\{X_{k}\}_{k\in\mathbb{N}}$ of infinite length had been generated.
@@
$\tau^{0}_{\epsilon,\delta}$ $=\min\left\{t\in\mathbb{N}:\,\mathbb{P}_{0}(|\mu(t)-\mu|>\epsilon)\leq\delta\right\}$
$\sim\min\left\{t\in\mathbb{N}:\,2\left(1-\Phi\left(\frac{\epsilon\sqrt{|M(t)|}}{v_{0}(t)}\right)\right)\leq\delta\right\},$ (4.4)
@@
$\quad=\mathbb{P}_{m(\tau^{0}_{\epsilon,\delta})}\left(\sqrt{|M(\tau^{0}_{\epsilon,\delta})|}\frac{|\mu_{\star}(\tau^{0}_{\epsilon,\delta})-\mu|}{v_{0}(\tau^{0}_{\epsilon,\delta})}>\sqrt{|M(\tau^{0}_{\epsilon,\delta})|}\frac{\epsilon}{v_{0}(\tau^{0}_{\epsilon,\delta})}\right)$
$\quad\approx 2\left(1-\Phi(\epsilon\sqrt{|M(\tau^{0}_{\epsilon,\delta})|}/v_{0}(\tau^{0}_{\epsilon,\delta}))\right),$ (4.5)
provided that the value of the stopping batch $\tau_{\epsilon,\delta}^{0}$ is luckily far away. We stress that the last approximation in (4.5) is valid because the regenerated batch $\{X_{k}^{\star}\}_{k\in M(\tau)}$ is conditionally independent of the stopping batch $\tau_{\epsilon,\delta}^{0}$ on the stopping-time $\sigma$-field $\mathcal{F}_{m(\tau_{\epsilon,\delta}^{0})}$.
@@ 4.2 The stopping rule based on implementable batch variances
As is now obvious from its expression, the crucial drawback of the approximate stopping rule (4.5) is the unavailability of the theoretical batch standard deviation $\{v_{0}(t)\}_{t\in\mathbb{N}}$. In what follows, we aim to replace the theoretical batch standard deviation in the approximate criterion (4.4) with a more implementable sequence $\{v(t)\}_{t\in\mathbb{N}}$ and make the resulting criterion more precise for a better application of the following martingale central limit theorem [P2012, Chapter VIII.1], which is a direct consequence of Theorem 4.2 owning to the continuous mapping theorem.
@@
In the literature, a variety of sufficient conditions have been provided in different forms for the weak convergence (4.6) to hold true, which we do not pursue as those are not directly relevant in our context. For detail, we refer the reader to, for instance, [ish, Theorem 2.5]. Let us stress that the condition $v^{2}(t)/v^{2}_{0}(t)\stackrel{{\scriptstyle\mathbb{P}_{0}}}{{\to}}1$ implies that the sequence $\{v(t)\}_{t\in\mathbb{N}}$ does not vanish in the limit, due to Assumption 4.1. From a practical point of view, this is a reasonable property, since the sample $\{X_{k}\}_{k\in\mathbb{N}}$ itself (not the sample average) satisfying $v^{2}(t)\to 0$ would otherwise necessarily converge to the constant $\mu$ of interest, for which a stopping rule would no longer be of particular use.
@@
$\tau_{\epsilon,\delta}:=\min\left\{t\in\mathbb{N}:\,2\left(1-\Phi\left(\frac{\epsilon\sqrt{|M(t)|}}{v(t)+a(t)}\right)\right)\leq\delta\right\},$ (4.7)
@@
Let us emphasize that the sequence $\{v(t)\}_{t\in\mathbb{N}}$ in (4.7) cannot be built on the regenerated sequence $\{X_{k}^{\star}\}_{k\in M(\tau_{\epsilon,\delta})}$, because $\tau_{\epsilon,\delta}$ needs to be an $(\mathcal{F}_{m(t)})_{t\in\mathbb{N}_{0}}$-stopping time, generated by the original sequence $\{X_{k}\}_{k\in\mathbb{N}}$ alone. On the basis of the stopping time (4.7), we construct a stopping rule as follows.
@@
In a similar manner to [YH1965], the “asymptotic” relevance of the criterion (4.7) can be justified in a way that the precision $\epsilon$ and the error probability $\delta$ are adjusted after the sequence $\{X_{k}\}_{k\in\mathbb{N}}$ of infinite length (which is certainly impossible to keep in hand) has been generated. We note that the almost sure convergence $v^{2}(t)/v^{2}_{0}(t)\to 1$ can be weakened to the weak convergence $\{v^{2}(\lfloor t/\epsilon\rfloor)/v^{2}_{0}(\lfloor t/\epsilon\rfloor):\,t\in(0,+\infty)\}\to 1$ in $\mathbb{D}((0,+\infty);\mathbb{R})$ as $\epsilon\to 0+$ for a version of Proposition 4.5 based on the functional weak law of large numbers, while it cannot be weakened further to $v^{2}(t)/v^{2}_{0}(t)\stackrel{{\scriptstyle\mathbb{P}_{0}}}{{\to}}1$ since then the corresponding version of Part (ii) would not hold anymore (see, for instance, [gut2009]).
@@
In relation to Proposition 4.5 (ii), we stress that $\mathbb{P}_{0}(|\mu(\tau_{\epsilon,\delta})-\mu|>\epsilon)\not\to\delta$ in general, as the random variable $\mu(\tau_{\epsilon,\delta})$ is not identical in law to the empirical batch mean $\mu(t)$ at a usual deterministic batch $t$ in the presence of the stopping time $\tau_{\epsilon,\delta}$. Conversely, the conditional standard deviation $s(\tau_{\epsilon,\delta})$ at the stopping batch $\tau_{\epsilon,\delta}$ cannot be replaced by its regenerated version, that is, $v_{\star}^{2}(\tau_{\epsilon,\delta})=|M(\tau_{\epsilon,\delta})|^{-1}\sum_{k\in M(\tau_{\epsilon,\delta})}{\rm Var}_{k-1}^{\star}(X^{\star}_{k})$, since the regenerated batch $\{X^{\star}_{k}\}_{k\in M(\tau)}$ is conditionally independent of the stopping batch $\tau_{\epsilon,\delta}$ on the stopping-time $\sigma$-field $\mathcal{F}_{m(\tau_{\epsilon,\delta})}$.
With Proposition 4.5 being presented, of practical interest is the non-asymptotic regime [BHST, FLYA2012] in which one generates a segment $\{X_{k}\}_{k\in M(t)}$ of the sequence, one batch at a time forward until, and only until, the criterion (4.7) is satisfied, with precision $\epsilon$ and error probability $\delta$ fixed.
@@ 4.2.1 Conditional batch variances
and $v_{1}(t):=(v_{1}^{2}(t))^{1/2}$ for $t\in\mathbb{N}$, where we have applied the base assumption (2.1) for the evaluation of the conditional variance. The conditional batch variance here is expected to behave in a somewhat similar manner to the theoretical one in the sense that it is unbiased for the theoretical batch variance (4.1), that is, $\mathbb{E}_{0}[v_{1}^{2}(t)]=v^{2}_{0}(t)$ for all $t\in\mathbb{N}$. Based on the availability of the sample $\{X_{k}\}_{k\in M(t)}$ alone, the conditional batch variance and standard deviation (4.8) are, in principle, more implementable than the theoretical ones (4.1), whereas the reality is that neither is generally implementable, with a few exceptions (such as the one examined in Section 5.1, where the conditional one (4.8) happens to be available in an implementable form).
@@
$\tau^{a}_{\epsilon,\delta}:=\min\left\{t\in\mathbb{N}:\,2\left(1-\Phi\left(\frac{\epsilon\sqrt{|M(t)|}}{v_{1}(t)+a(t)}\right)\right)\leq\delta\right\},$ (4.9)
@@ 4.2.2 Empirical batch variances
$\tau^{b}_{\epsilon,\delta}:=\min\left\{t\in\mathbb{N}:\,2\left(1-\Phi\left(\frac{\epsilon\sqrt{|M(t)|}}{v_{2}(t)+a(t)}\right)\right)\leq\delta\right\},$ (4.11)
@@ 5 Numerical illustrations
In accordance with [BHST], we consider a stopping rule with its reliability $R_{\epsilon,\delta}(\omega)$ strictly less than $1$ unreliable. In addition, the effectiveness of a stopping rule improves as its complexity $\text{CM}_{\epsilon,\delta}(\omega)$ decreases at the specified precision and error probability. Throughout the following experiments, we fix $n=5\times 10^{3}$ to ensure enough realizations. We note that the parameter $\ell$ in the complexity $\text{CM}_{\epsilon,\delta}(\omega)$ controls how severely we penalize large stopping times and is suppressed from its notation because we fix $\ell=5$ throughout.
@@
In both examples below (Sections 5.1 and 5.2), the theoretical variance $v_{0}^{2}(t)$ converges to a positive constant, so that $|M(t+1)|\sim|M(t)|$ to satisfy the technical condition $v_{0}^{2}(t+1)/|M(t+1)|\sim v_{0}^{2}(t)/|M(t)|$ stated in Assumption 4.1. Accordingly, we set the batch sizes as $M(t)=\{(t-1)^{5}+1,\cdots,t^{5}\}$ so that $|M(t)|\propto t^{4}$, striking a practical balance. Slower growth (for instance, $|M(t)|\propto t^{2}$) reduces excess sampling but requires more frequent checks, whereas faster growth (for instance, $|M(t)|\propto t^{6}$) decreases the number of checks but risks overshooting the required sample size.
@@ 5.1 ARCH(1)
We begin our numerical illustrations with an ARCH(1) model, which provides an ideal test case for our stopping rules due to its non-trivial dependence structure and heavy-tailed behavior, as well as the straightforward implementation of the theoretical and conditional batch variances. This setting poses natural challenges for sequential sampling methods, as the underlying dynamics combine temporal dependence through the ARCH structure with potentially extreme observations from the $t$-distribution.
Now, consider an ARCH(1) sequence $\{X_{k}\}_{k\in\mathbb{N}}$, which is generated recursively by $X_{k}=(\beta+\alpha X_{k-1}^{2})^{1/2}V_{k}$, with $X_{0}=0$, $\alpha\in(0,1)$ and $\beta\in(0,+\infty)$. Here, $\{V_{k}\}_{k\in\mathbb{N}}$ denotes a sequence of iid random variables following a scaled $t$-distribution with $\nu$ degrees of freedom ($\nu\in\{5,6,\cdots\}$), normalized to have unit variance, with its probability density function given as
$\frac{d}{dx}\mathbb{P}(V_{1}\leq x)=\frac{\Gamma((\nu+1)/2)}{\sqrt{(\nu-2)\pi}\Gamma(\nu/2)}\left(1+\frac{x^{2}}{\nu-2}\right)^{-(\nu+1)/2},\quad x\in\mathbb{R}.$
@@
In what follows, we fix $\alpha=0.03$, $\beta=0.3$, and $\nu=6$ degrees of freedom. As we observe from Figure 2, after some initial fluctuations due to smaller batch sizes, the theoretical, conditional, and empirical variances exhibit similar behavior, with all those measures tracking each other closely. This validates the use of the empirical variance $v_{2}^{2}(t)$ in the stopping rule, as it succeeds in approximating the theoretical variance, not to mention the conditional variance, in the long run.
Figure 2: Typical realization of theoretical batch variance $v_{0}^{2}(t)$ (solid blue), conditional batch variance $v_{1}^{2}(t)$ (dashed red), and empirical batch variance $v_{2}^{2}(t)$ (dotted green) of the ARCH(1) process under consideration. The inset displays the absolute estimation errors $|v_{1}^{2}(t)-v_{0}^{2}(t)|$ (dashed red) and $|v_{2}^{2}(t)-v_{0}^{2}(t)|$ (dotted green) on a logarithmic scale, illustrating the convergence rates to the theoretical variance.
@@
We implement the empirical variance stopping rule $\tau^{b}_{\epsilon,\delta}$ based on the empirical batch variance $v_{2}^{2}(t)$ as of (5.1), as it can be readily available unlike the theoretical and conditional variances. Here, we set $a(t)=1/t$, which helps to compensate the severe underestimation of the empirical variance $v_{2}^{2}(t)$ during early batches, such as $v_{2}^{2}(1)\approx 0$ observed in Figure 2. The choice of $a(t)=1/t$ satisfies its theoretical requirement of monotonic vanishing and reduces the risk of premature stopping.
@@
We assess its performance by examining reliability and computational complexity ((iv)) across various precision ($\epsilon$) and error probability ($\delta$) values. For each $(\epsilon_{k},\delta_{k})$ in the specified grid, we perform independent simulations to compute $\tau^{b}_{\epsilon_{k},\delta_{k}}(\omega_{j})$, generate new batches for the regenerated mean $\mu_{\star}(\tau_{\epsilon_{k},\delta_{k}})(\omega_{j})$, and evaluate success through $\mathbbm{1}(|\mu_{\star}(\tau_{\epsilon_{k},\delta_{k}})(\omega_{j})-\mu|\leq\epsilon_{k})$.
Figure 3: The left column ((a) and (c)) represents the reliability $R_{\epsilon,\delta}$, while the right column ((b) and (d)) shows the complexity ${\rm CM}_{\epsilon,\delta}$ on a $\log_{10}$ scale. The top row ((a) and (b)) corresponds to the empirical-variance stopping rule ($\tau^{b}_{\epsilon,\delta}$), whereas the bottom row ((c) and (d)) corresponds to the conditional-variance stopping rule ($\tau^{a}_{\epsilon,\delta}$).
Table 1: Performance metrics comparison between the conditional-variance ($\tau^{a}_{\epsilon,\delta}$) and empirical-variance ($\tau^{b}_{\epsilon,\delta}$) stopping rules.
Metric Mean Min Max
Reliability $R_{\epsilon,\delta}$
     empirical (Figure 3 (a)) 1.0165 1.0000 1.1111
     conditional (Figure 3 (c)) 1.0191 1.0001 1.1111
Complexity $\rm CM_{\epsilon,\delta}$
     empirical (Figure 3 (b)) 1,965,883.60 243.00 24,300,000.00
     conditional (Figure 3 (d)) 3,339,877.00 1,024.00 45,435,424.00
In Figure 3 and Table 1 above, we present our comprehensive numerical findings on the performance of the proposed stopping rules. The plots of the reliability $R_{\epsilon,\delta}$ (Figure 3 (a) and (c)) demonstrate that our stopping rules consistently achieve the reliability staying strictly greater than $1$ across all parameter combinations. As confirmed in Table 1, the reliability ranges from 1.000 to 1.111, satisfying the reliability criterion set up in [BHST]. Notably, the conditional-variance stopping rule $\tau^{a}_{\epsilon,\delta}$ exhibits a slightly higher mean (1.0191) compared to the empirical-variance stopping rule $\tau^{b}_{\epsilon,\delta}$ (1.0165), highlighting the conservative nature of using the true model parameters. The reliability increases systematically with larger values of $\epsilon$ and $\delta$, reflecting more conservative behavior in less demanding scenarios while maintaining reliability even in the most challenging case with $\epsilon=\delta=10^{-3}$.
The heat maps of the complexity ${\rm CM}_{\epsilon,\delta}$ (Figure 3 (b) and (d)) reveal a pattern of computational cost across the parameter space. The logarithmic scale clearly shows how the complexity $\rm CM_{\epsilon,\delta}$ increases as both $\epsilon$ and $\delta$ decrease. The computational burden varies substantially, as detailed in Table 1, ranging from $2.43\times 10^{2}$ for the empirical-variance stopping rule with larger values of $\epsilon$ and $\delta$ to $4.54\times 10^{7}$ for the conditional-variance stopping rule with $\epsilon=\delta=10^{-3}$. Directly comparing the methods, the empirical-variance stopping rule requires approximately $2.0\times 10^{6}$ samples on average versus $3.3\times 10^{6}$ for the conditional-variance stopping rule, demonstrating that the empirical approach achieves significant computational efficiency without violating the reliability requirement. This scaling aligns with our theoretical predictions and provides practical guidance for computational requirements.
Let us stress that smooth transitions in both reliability and complexity across the parameter space indicate the stability and robustness of the proposed stopping rules. The consistent achievement of the requirement ($R_{\epsilon,\delta}>1$), coupled with theoretically predicted complexity scaling, provides strong empirical validation of our methodology, particularly notable given the challenging features of temporal dependence and heavy-tailed distributions in the ARCH(1) setting.